# Light Control of Triplet Pairing in Correlated Electrons with Mixed-Sign Interactions

### v2

@@ Abstract
Abstract: Spin-triplet superconductivity is a key platform for topological quantum computing, yet its experimental realization and control in solid-state materials remain a significant challenge. For this purpose, we propose an ultrafast optical strategy to manipulate spin-triplet superconductivity by leveraging $p$-wave pairing instabilities in the extended Hubbard model, a framework applicable to transition-metal oxides. Utilizing Floquet engineering, we demonstrate that transient flipping of the effective spin-exchange interaction can enhance $p$-wave pairing correlations under linearly polarized optical pulses. Furthermore, we reveal that this emergent spin-triplet pairing in strongly correlated systems can be selectively switched by an orthogonal optical pulse. This work provides a new pathway for stabilizing and controlling spin-triplet superconductivity in correlated materials.
@@ I Introduction
Spin-triplet superconductivity is widely recognized as a promising avenue for topological quantum computing [1, 2], offering a robust platform for hosting non-Abelian excitations and Majorana fermions [3, 4, 5]. However, despite its theoretical appeal, experimental realization remains an ongoing challenge. Sr2RuO4 was long considered the leading candidate, particularly for a chiral $p+ip$ pairing state [6, 7, 8, 9, 10, 11], but recent experimental findings have cast significant doubt on its pairing symmetry [12, 13, 14]. Other materials, such as UTe2[15, 16], UPt3[17], and K2Cr3As3[18] have been proposed as potential candidates for spin-triplet superconductors, yet unambiguous experimental validation of spin-triplet pairing remain elusive.
Beyond specific materials, theoretical studies and numerical simulations have predicted the emergence of short-range or quasi-long-range spin-triplet pairing correlations in several strongly correlated models with mixed-sign interactions. Notably, $p$-wave superconductivity has been identified in the one-dimensional (1D) extended Hubbard model (EHM), where a repulsive on-site interaction $U$ coexists with an attractive nearest-neighbor interaction $V$, favoring nodal pairing symmetry. Under specific parameter regimes and doping conditions, this model exhibits dominant spin-triplet pairing instability [19, 20, 21, 22, 23, 24]. Extending this model to two-dimensional (2D) further reveals a strong spin-triplet pairing instability near quarter filling [25, 26, 27, 28, 29, 30, 31]. While initially considered an artificial toy model, mixed-sign interactions have been experimentally identified in strongly correlated cuprates in quasi-1D thin films and single crystals [32, 33, 34, 35]. In these systems, electron-phonon coupling is believed to mediate attractive interactions between nearest-neighbor electrons that remain weaker than the on-site repulsion, leading to a scenario effectively described by the mixed-sign interaction model [36, 37, 38, 39, 40]. Although additional experimental validation is needed to quantify these interactions, such a mechanism could be widely realized in Ruddlesden-Popper transition-metal oxides, providing a promising route for stabilizing spin-triplet pairing.
@@
Another challenge in harnessing spin-triplet superconductivity is achieving efficient and selective control over its order parameter. Recent time-dependent mean-field studies have proposed sophisticated strategies for manipulating multiple $p$-wave order parameters [41], and inducing transitions between spin-singlet and triplet states [42]. In mean-field approximations, where quantum fluctuations are neglected, these approaches typically require carefully timed sequences of pump pulses with varying polarizations. In strongly correlated systems, light-induced superconductivity may be more directly accessible due to the interplay of competing orders and quantum fluctuations. Experimental evidences from cuprates and molecular crystals suggest that nonequilibrium optical excitation can dynamically alter superconducting states in correlated materials [43, 44, 45, 46, 47, 48].
Building on these developments, we propose a dynamical strategy to engineer spin-triplet superconductivity in strongly correlated systems with mixed-sign interactions. Specifically, we leverage the Floquet engineering of magnetic exchange interactions [49, 50] to induce a transient sign reversal of the exchange interaction along the pump polarization, when the pump laser is tuned near resonance [see Fig. 1]. Using time-dependent exact diagonalization simulations, we demonstrate the feasibility of selectively enhancing existing spin-triplet pairing correlations on ultrafast timescales. A subsequent pulse with an orthogonal polarization can further manipulate the order parameter, enabling controlled switching between pairing symmetries. This approach paves the way for designing and dynamically controlling spin-triplet superconductivity in strongly correlated materials.
@@ II.1 Light-Enhanced Spin-Triplet Pairing Correlations
To investigate spin-triplet pairing instabilities in strongly correlated systems, we focus on a quarter-filled (i.e., 50% hole-doped) EHM on a square lattice, which serves as a prototypical model for transition-metal oxides. Following Eq. (6) in Methods, the EHM incorporates a repulsive on-site interaction $U$ and an attractive nearest-neighbor interaction $V$. Previous studies have demonstrated that such a model supports strong spin-triplet pairing instabilities [19, 20, 21, 22, 23, 25, 26, 27, 28], confirmed through many-body numerical simulations [24, 29, 30, 31]. Here, we adopt $U=8t_{h}$ and $V=-t_{h}$ as model parameters in Eq. (6), consistent with experimental estimates for cuprates [32, 33, 34].
The spin-triplet pairing instabilities are characterized by the three-fold $p$-wave pairing operators for each spatial symmetry. For example, the $p_{x}$-wave pairing operator is defined as:
@@
where $c_{\mathbf{i},\sigma}$ represents the annihilation operator for an electron with spin $\sigma$ at site $\mathbf{i}$. Analogous operators for the $p_{y}$, $p_{x}\pm p_{y}$, and $p_{x}\pm ip_{y}$ pairing channels are defined in the Supplementary Note I. For any many-body state $|\psi\rangle$, the $p$-wave pairing correlations are computed as:
@@
where $\alpha$ labels the pairing symmetry and the pairing operator is averaged over the triplet states $\Delta^{\alpha}_{\mathbf{i}}=(\Delta^{\alpha}_{\mathbf{i},1}+\Delta^{\alpha}_{\mathbf{i},0}+\Delta^{\alpha}_{\mathbf{i},-1})/\sqrt{3}$. These pairing correlations can be represented as vectors on the Bloch sphere (see Fig. 1b), providing a geometric perspective on the interplay between different pairing symmetries. In this work, we focus on analyzing pairing correlations within finite clusters rather than long-range order, the latter of which remains challenging to establish in 2D systems due to strong quantum fluctuations.
@@
At equilibrium, the $p$-wave pairing correlations, which are isotropic across all six orientations, exhibit a pronounced hump for the interaction parameters adopted in this study ($U=8t_{h}$ and $V=-t_{h}$), signaling strong spin-triplet pairing instabilities [29]. Starting from this equilibrium state, we investigate the nonequilibrium dynamics of $p$-wave pairing correlations induced by an external optical excitation. The time-dependent wavefunction $|\psi(t)\rangle$ is obtained by evolving the system under a time-dependent Hamiltonian, where the external vector potential $\mathbf{A}(t)$ is incorporated via the Peierls substitution (see Methods). The instantaneous pairing correlation $\Phi_{\alpha}(t)$ is then calculated by substituting $|\psi(t)\rangle$ into Eq. (2). For the single-pump setup considered in this section, the external light pulse is usually modeled as a Gaussian envelope:
@@
Without loss of generality, we focus on a single $x$-polarized laser pulse, $\hat{\textbf{e}}_{\rm pol}=\hat{x}$, to explore the anisotropic response of pairing correlations. In typical pump-probe experiments, phase coherence between the pump and probe lasers is often not strictly maintained, leading to phase-averaged measurements that suppress phase-locked oscillations. To account for this, we evaluate the phase-averaged pairing correlation as $\sum_{n}\Phi_{\alpha}(t;\phi=\phi_{n})/n$, with a detailed analysis provided in the Supplementary Note II. For clarity, all figures in the main text present results with $\phi=0$, as the phase slightly influences the amplitude of $\Phi_{\alpha}(t)$ without qualitatively altering its overall temporal evolution.
@@
The dynamics of spin-triplet pairing correlations under an $x$-polarized laser pulse with pump frequency $\Omega=9.2t_{h}$ are presented in Figs. 2a-c. A prominent feature of this response is the pronounced oscillatory behavior observed in $\Phi_{x}(t)$ and $\Phi_{y}(t)$, which exhibit a striking anti-phase relationship both during and after the pump pulse. This anti-phase behavior reflects an intrinsic competition between these orthogonal pairing channels. While increasing the pump strength enhances the oscillation amplitude, the underlying anti-phase structure remains unchanged, indicating the robustness of this dynamical competition. As a result, the pairing channel aligned with the pump field (e.g., $p_{x}$) is transiently enhanced at specific oscillation phases, such as near $t=10t_{h}^{-1}$, at the cost of significantly suppressing its orthogonal counterpart (e.g., $p_{y}$).
@@
In contrast, other orthogonal pairs of pairing channels, such as $\Phi_{x\pm y}(t)$, exhibit a different dynamical response [see Fig. 2c]. These channels undergo an abrupt suppression during the pump pulse and show minimal oscillatory behavior. Instead, their post-pump evolution reflects an averaged behavior of $\Phi_{x}(t)$ and $\Phi_{y}(t)$, suggesting a net suppression of pairing correlations, likely driven by photo-induced carrier screening effects [51]. The maximum enhancement of $\Phi_{x}(t)$ follows a non-monotonic dependence on pump strength, dictated by the competing effects of light-enhanced $p_{x}$ pairing and the overall suppression of pairing correlations.
@@
To further elucidate the anisotropic response of laser-induced $p$-wave pairing correlations, we map the pairing correlations onto the Bloch sphere at four representative time points: $t=-8t_{h}^{-1}$, $2t_{h}^{-1}$, 0 and $7t_{h}^{-1}$ [see Fig. 2d]. At equilibrium, pairing correlations are isotropic without any symmetry breaking. However, once a linearly polarized pump is applied, this symmetry is broken, causing the Bloch sphere to shift preferentially toward the $p_{x}$ direction while suppressing the $p_{y}$ component. This asymmetry becomes more pronounced with higher pump fluence, indicating a stronger laser-induced anisotropy. After the pump, the Bloch sphere undergoes continuous oscillations along the horizontal axis, yet $\Phi_{x}(t)$ consistently dominates over $\Phi_{y}(t)$. Notably, the center of the Bloch sphere does not cross the origin, as shown by the directional arrows in Fig. 2d, reflecting sustained $p_{x}$ channel dominance triggered by the pump. The first maximal imbalance occurs near $t\sim\sigma=2.5t_{h}^{-1}$, with slight variations depending on the pump amplitude. Since the system is closed and lacks dissipation, the undamped laser energy drives continuous Bloch sphere oscillations over extended timescales, ensuring the persistence of the $p_{x}$ dominance and enhancement of spin-triplet pairing.
@@ II.2 Floquet Engineering of Anisotropic Spin Interactions
Figure 4: Floquet-engineered anisotropic spin exchange. a Schematic depiction of the spin exchange process in a Floquet system. The effective superexchange interaction $J$ originates from second-order virtual hopping processes that involve multiple intermediate states, with their energies set by the equilibrium intermediate-state energy $E_{\text{int}}$ and integer multiples of the pump frequency $\Omega$. b Effective spin exchange interaction $J_{\alpha}^{\rm eff}$ at a transient state, calculated using the Floquet theory, as a function of the pump frequency $\Omega$. The red and blue regions indicate ferromagnetism (FM) and antiferromagnetism (AFM) exchange interactions, respectively. The solid lines correspond to calculations assuming homogeneous electron occupations, while the dashed lines represent redshifted and blueshifted $J_{\alpha}^{\rm eff}$ due to occupation imbalances, as illustrated in c. The single- and two-photon resonances are determined by the variation of FM regions. c Schematic illustration of various intermediate state configurations with distinct electron occupations and their corresponding intermediate-state energies $E_{\text{int}}$.
The observed frequency-dependent enhancement of spin-triplet pairing correlations can be attributed to Floquet engineering of effective spin-exchange interactions. In the presence of a periodic pump field, the transient system can be approximated by an effective Floquet Hamiltonian, treating the laser field as periodic [52, 53, 54]. Within this framework, the monochromatic light field couples to the electronic states, creating a series of photon-dressed states with energy levels spaced by multiples of the pump frequency $\Omega$[55]. In strongly correlated systems, such photon dressing significantly modifies the intermediate states involved in spin superexchange process [49, 50]. As illustrated in Fig. 4a, an intermediate state with double occupancy transforms into photon-dressed states under periodic driving, effectively renormalizing the spin superexchange interactions. Consequently, the Floquet-engineered spin exchange interaction becomes [49, 50]:
@@
where $\mathcal{J}_{m}(x)$ is the $m$th-order Bessel function of the first kind, describing contributions from processes involving $m$ photons with energy $\Omega$. Here, $E_{\rm int}$ represents the energy cost of creating a doubly occupied intermediate state relative to the singlet ground state. Crucially, the effective exchange interaction $J_{\alpha}^{\mathrm{eff}}$ depends on the projection $A_{0}\cdot\hat{\mathbf{e}}_{\alpha}$ of the pump field along the exchange pathway, allowing dynamic control over both the magnitude and the sign of the interaction by tuning the pump polarization, amplitude, and frequency [56].
@@
In contrast to the Hubbard model, where the intermediate-state energy is determined by $U$ alone [49, 50], the EHM with on-site $U$ and nearest-neighbor $V$ results in an intermediate-state energy $E_{\mathrm{int}}\approx U-V$, setting the single-photon resonance at $\Omega\approx 9t_{h}$ in Eq. (4). As shown in Fig. 4b, the calculated spin exchange integral $J_{x}^{\rm eff}$ for $A_{0}=0.4$ exhibits strong pump frequency dependence (see Fig. 4b). Away from resonance, the Floquet-engineered effective $J_{x}^{\rm eff}$ recovers its equilibrium antiferromagnetic (AFM) exchange interaction $J\sim 4t_{h}^{2}/E_{\mathrm{int}}\sim 0.5t_{h}$. However, as $\Omega$ approaches resonance with $E_{\mathrm{int}}$, $J_{x}^{\rm eff}$ undergoes significant renormalization. Notably, when $\Omega$ slightly exceeds $E_{\rm int}$, the exchange interaction undergoes a sign reversal, transitioning from AFM to ferromagnetic (FM) (red-shaded region). This sign flip is crucial because AFM exchange favors spin-singlet pairing, while FM exchange supports spin-triplet pairing, thereby explaining the enhancement of $\Phi_{x}$ for $\Omega\gtrsim E_{\mathrm{int}}$. Moreover, the Floquet effect is polarization-selective: the exchange interaction along the $y$ direction ($J_{y}^{\rm eff}$) remains largely unaffected due to its orthogonality with the pump polarization, leading to the observed anisotropic pairing correlations.
@@
The resonance condition for Floquet-induced FM exchange is not confined to the narrow frequency window denoted by the red-shaded regime in Fig. 4b, but extends across a broader range of pump frequencies. This broadening arises from two primary effects. First, the above simplified Floquet analysis, which is based on atomic models, requires adjustments to incorporate band dispersion and quantum fluctuations in many-body systems. Second, the intermediate-state energy $E_{\mathrm{int}}$, which determines $J_{x}^{\rm eff}$ in Eq. (4), is not a fixed quantity in the doped EHM but varies across lattice sites in a given Fock state, depending on the local occupation configurations. As illustrated in the upper panels of Fig. 4c, configurations with isolated spin pairs or uniform neighbor densities yield $E_{\rm int}=U-V$, corresponding to a strong resonance near $\Omega\sim U-V=9t_{h}$ (solid lines in Fig. 3b). However, non-uniform occupations introduce a distribution of $E_{\rm int}$ values spanning from U4VU-4V to $U+2V$. At an average filling of $\langle n\rangle=0.5$, this variation extends the resonance window from $U+0.5V=7.5t_{h}$ to $U-2.5V=10.5t_{h}$, as indicated by dashed lines in Fig. 4b. This broadened resonance range aligns with the observed light-enhanced $p_{x}$ pairing instabilities in Fig. 3, reflecting the role of Floquet engineering in tuning pairing correlations.

### v3

@@ Abstract
Abstract: Spin-triplet superconductivity is a key platform for topological quantum computing, yet its experimental realization and control in solid-state materials remain a significant challenge. For this purpose, we propose an ultrafast optical strategy to manipulate spin-triplet superconductivity by leveraging $p$-wave pairing instabilities in the extended Hubbard model, a framework applicable to transition-metal oxides. Utilizing Floquet engineering, we demonstrate that transient flipping of the effective spin-exchange interaction can enhance $p$-wave pairing correlations under linearly polarized optical pulses. Furthermore, we reveal that this emergent spin-triplet pairing in strongly correlated systems can be selectively switched by an orthogonal optical pulse. This work provides a pathway for stabilizing and controlling spin-triplet superconductivity in correlated materials.
@@ I Introduction
Spin-triplet superconductivity is widely recognized as a promising avenue for topological quantum computing [sarma2015majorana, nayak2008non], offering a robust platform for hosting non-Abelian excitations [read2000paired, ivanov2001non, kitaev2001unpaired]. However, despite its theoretical appeal, experimental realization remains an ongoing challenge. Sr2RuO4 was long considered the leading candidate, particularly for a chiral $p+ip$ pairing state [maeno1994superconductivity, rice1995sr2cuo4, baskaran1996why, mackenzie2003the, maeno2011evaluation, nelson2004odd], but recent experimental findings have cast significant doubt on its pairing symmetry [pustogow2019constraints, ishida2020reduction, chronister2021evidence]. Other materials, such as UTe2[ran2019nearly, ran2020enhancement], UPt3[tou1996odd], and K2Cr3As3[yang2021spin] have been proposed as potential candidates for spin-triplet superconductors, yet unambiguous experimental validation of spin-triplet pairing remain elusive.
Beyond specific materials, theoretical and numerical studies have revealed that spin-triplet pairing can naturally emerge in strongly correlated models with mixed-sign interactions. In particular, the one-dimensional (1D) extended Hubbard model (EHM), where a repulsive on-site interaction $U$ coexists with an attractive nearest-neighbor interaction $V$, has been shown to support nodal pairing symmetry. Under appropriate doping and parameter regimes, this model exhibits spin-triplet superconductivity [lin1986condensation, penc1994phase, lin1995phase, xiang2019doping, shinjo2019machine, qu2022spin]. Its extension to two dimensions (2D) further reveals robust spin-triplet pairing instability near quarter filling [onari2004phase, huang2013unconventional, nayak2018exotic, van2018extended, chen2023superconducting, cao2025dominant, farid2025lifshitz]. Although once treated as an artificial toy model, mixed-sign interactions have been experimentally identified in strongly correlated cuprates in quasi-1D thin films and single crystals [chen2021anomalously, padma2025beyond, scheie2025cooper, Li2024Doping]. In these systems, phonons are believed to mediate attractive interactions between nearest-neighbor electrons while preserving dominant on-site repulsion, leading to a scenario effectively described by the EHM [wang2021phonon, tang2023traces, feiguin2023effective, shen2024signatures, tang2024influence]. Although further experimental validation is needed to quantify these interactions, such mechanisms may be broadly realized in Ruddlesden-Popper transition-metal oxides due to structural similarities, providing a promising route toward stabilizing spin-triplet pairing beyond quasi-1D materials.
@@
Another challenge in harnessing spin-triplet superconductivity is achieving efficient and selective control over its order parameter. Recent time-dependent mean-field studies have proposed sophisticated strategies for manipulating multiple $p$-wave order parameters [claassen2019universal], and inducing transitions between spin-singlet and triplet states [gassner2024light]. In mean-field approximations, where quantum fluctuations are neglected, these approaches typically require carefully timed sequences of pump pulses with varying polarizations. In strongly correlated systems, light-induced superconductivity may be more directly accessible due to the interplay of competing orders and quantum fluctuations. Experimental evidences from cuprates and molecular crystals suggest that nonequilibrium optical excitation can dynamically alter superconducting states in correlated materials [fausti2011light, nicoletti2014optically, hu2014optically, mankowsky2014nonlinear, buzzi2020photomolecular, buzzi2021phase].
Building on these developments, we propose a dynamical strategy to engineer spin-triplet pairing in strongly correlated systems with mixed-sign interactions. Specifically, we leverage the Floquet engineering of magnetic exchange interactions [mentink2014ultrafast, mentink2015ultrafast] to analysis how a transient sign reversal of the exchange interaction along the pump polarization, when the pump laser is tuned near resonance [see Fig. 1]. Using time-dependent exact diagonalization simulations on a extended Hubbard model, we demonstrate the feasibility of selectively enhancing existing spin-triplet pairing correlations on ultrafast timescales. A subsequent pulse with an orthogonal polarization can further manipulate the order parameter, enabling controlled switching between pairing symmetries. This approach paves the way for designing and dynamically controlling spin-triplet superconductivity in strongly correlated materials.
@@ II.1 Light-Enhanced Spin-Triplet Pairing Correlations
To investigate spin-triplet pairing instabilities in strongly correlated systems, we focus on a quarter-filled (i.e., 50% hole-doped) EHM on a square lattice, which serves as a prototypical model for transition-metal oxides. Following Eq. (6) in Methods, the EHM incorporates a repulsive on-site interaction $U$ and an attractive nearest-neighbor interaction $V$. Previous studies have demonstrated that such a model supports strong spin-triplet pairing instabilities [lin1986condensation, penc1994phase, lin1995phase, xiang2019doping, shinjo2019machine, onari2004phase, huang2013unconventional, nayak2018exotic, van2018extended], confirmed through many-body numerical simulations [qu2022spin, chen2023superconducting, cao2025dominant, farid2025lifshitz]. Here, we adopt $U=8t_{h}$ and $V=-t_{h}$ as model parameters in Eq. (6), consistent with experimental estimates for cuprates [chen2021anomalously, padma2025beyond, scheie2025cooper].
The spin-triplet pairing instabilities are characterized by the three-fold pairing operators for each spatial symmetry. For example, the $p_{x}$-wave pairing operator is:
@@
where $c_{\mathbf{i},\sigma}$ represents the annihilation operator for an electron with spin $\sigma$ at site $\mathbf{i}$. Analogous operators for the $p_{y}$, $p_{x}\pm p_{y}$, and $p_{x}\pm ip_{y}$ pairing channels are defined in the Supplementary Note 1. For any many-body state $|\psi\rangle$, the $p$-wave pairing correlations are computed as:
@@
where $\alpha$ labels the pairing symmetry and the pairing operator is averaged over the triplet states $\Delta^{p_{\alpha}}_{\mathbf{i}}=(\Delta^{p_{\alpha}}_{\mathbf{i},1}+\Delta^{p_{\alpha}}_{\mathbf{i},0}+\Delta^{p_{\alpha}}_{\mathbf{i},-1})/\sqrt{3}$. These pairing correlations can be represented as vectors on the Bloch sphere (see Fig. 1b), providing a geometric perspective on the interplay between different pairing symmetries. In this work, we focus on analyzing short-range pairing correlations within finite clusters rather than long-range order, the latter of which remains challenging to establish in 2D systems due to strong quantum fluctuations.
@@
At equilibrium, the $p$-wave pairing correlations, which are isotropic across all six orientations, exhibit a pronounced hump for the interaction parameters adopted in this study ($U=8t_{h}$ and $V=-t_{h}$), signaling strong spin-triplet pairing instabilities [chen2023superconducting]. Starting from this equilibrium state, we investigate the nonequilibrium dynamics of $p$-wave pairing correlations induced by an external optical excitation. The time-dependent wavefunction $|\psi(t)\rangle$ is obtained by evolving the system under a time-dependent Hamiltonian, where the external vector potential $\mathbf{A}(t)$ is incorporated via the Peierls substitution (see Methods). The instantaneous pairing correlation $\Phi_{\alpha}(t)$ is then calculated by substituting $|\psi(t)\rangle$ into Eq. (2). For the single-pump setup considered in this section, the external light pulse is usually modeled as a Gaussian envelope:
@@
Without loss of generality, we focus on a single $x$-polarized laser pulse, $\hat{\textbf{e}}_{\rm pol}=\hat{x}$, to explore the anisotropic response of pairing correlations [polarizationSymmetry]. In typical pump-probe experiments, phase coherence between the pump and probe lasers is often not strictly maintained, leading to phase-averaged measurements that suppress phase-locked oscillations. To account for this, we evaluate the phase-averaged pairing correlation as $\sum_{n}\Phi_{\alpha}(t;\phi=\phi_{n})/n$, with a detailed analysis provided in the Supplementary Note 2. For clarity, all figures in the main text present results with $\phi=0$, as the phase slightly influences the amplitude of $\Phi_{\alpha}(t)$ without qualitatively altering its overall temporal evolution.
@@
The dynamics of spin-triplet pairing correlations under an $x$-polarized laser pulse with $\Omega=9.2t_{h}$ are presented in Figs. 2a-c. A prominent feature of this response is the pronounced oscillatory behavior observed in $\Phi_{x}(t)$ and $\Phi_{y}(t)$, which exhibit a striking anti-phase relationship both during and after the pump pulse. This anti-phase behavior reflects an intrinsic competition between these orthogonal pairing channels. While increasing the pump strength enhances the oscillation amplitude, the underlying anti-phase structure remains unchanged, indicating the robustness of this dynamical competition. As a result, the pairing channel aligned with the pump field (e.g., $p_{x}$) is transiently enhanced at specific oscillation phases, such as near $t=10t_{h}^{-1}$, at the cost of significantly suppressing its orthogonal counterpart (e.g., $p_{y}$).
@@
In contrast, other orthogonal pairs of pairing channels, such as $\Phi_{x\pm y}(t)$, exhibit a different dynamical response [see Fig. 2c]. These channels undergo an abrupt suppression during the pump pulse and show minimal oscillatory behavior. Instead, their post-pump evolution reflects an averaged behavior of $\Phi_{x}(t)$ and $\Phi_{y}(t)$, suggesting a net suppression of pairing correlations, likely driven by photo-induced carrier screening effects [wang2021fluctuating]. The maximum enhancement of $\Phi_{x}(t)$ follows a non-monotonic dependence on pump strength, dictated by the competing effects of light-enhanced $p_{x}$ pairing and the overall suppression of pairing correlations.
@@
To further elucidate the anisotropic response of laser-induced $p$-wave pairing correlations, we map the pairing correlations onto the Bloch sphere at four representative time points: $t=-8t_{h}^{-1}$ , $2t_{h}^{-1}$, 0 and $7t_{h}^{-1}$ [see Fig. 2d]. At equilibrium, pairing correlations are isotropic without any symmetry breaking. However, once a linearly polarized pump is applied, this symmetry is broken, causing the Bloch sphere to shift preferentially toward the $p_{x}$ direction while suppressing the $p_{y}$ component. This asymmetry becomes more pronounced with higher pump fluence, indicating a stronger laser-induced anisotropy. After the pump, the Bloch sphere undergoes continuous oscillations along the horizontal axis, yet $\Phi_{x}(t)$ consistently dominates over $\Phi_{y}(t)$. Notably, the center of the Bloch sphere does not cross the origin, as shown by the directional arrows in Fig. 2d, reflecting sustained $p_{x}$ channel dominance triggered by the pump. The first maximal imbalance occurs near $t\sim\sigma=2.5t_{h}^{-1}$, with slight variations depending on the pump amplitude. Since the correlated system is closed and lacks dissipation, the undamped laser energy drives continuous Bloch sphere oscillations over extended timescales lie within the pre-thermal plateau [aoki2014nonequilibrium, de2021colloquium], ensuring the persistence of the $p_{x}$ dominance and enhancement of spin-triplet pairing after the pump pulse ends [kalthoff2019floquet, sentef2020quantum].
@@ II.2 Floquet Engineering of Anisotropic Spin Interactions
Figure 4: Floquet-engineered anisotropic spin exchange. a Schematic depiction of the spin exchange process in a Floquet system. The effective superexchange interaction $J$ originates from second-order virtual hopping processes that involve multiple intermediate states, with their energies set by the equilibrium intermediate-state energy $E_{\text{int}}$ and integer multiples of the pump frequency $\Omega$. b Effective spin exchange interaction $J_{\alpha}^{\rm eff}$ at a transient state, calculated using the Floquet theory, as a function of the pump frequency $\Omega$. The red and blue regions indicate ferromagnetism (FM) and antiferromagnetism (AFM) exchange interactions, respectively. The solid lines correspond to calculations assuming homogeneous electron occupations, while the translucent extended $J_{\alpha}^{\rm eff}$ resonance area due to occupation imbalances, as illustrated in c. The single- and two-photon resonances are determined by the variation of FM regions. c Schematic illustration of various intermediate state configurations with distinct electron occupations and their corresponding intermediate-state energies $E_{\text{int}}$.
The observed frequency-dependent enhancement of spin-triplet pairing correlations in the dynamical simulation of the EHM model can be attributed to Floquet engineering of effective spin-exchange interactions. In the presence of a periodic pump field, the transient system can be approximated by an effective Floquet Hamiltonian, treating the laser field as periodic [oka2019floquet, rudner2020band, de2021colloquium]. Within this framework, the monochromatic light field couples to the electronic states, creating a series of photon-dressed states with energy levels spaced by multiples of the pump frequency $\Omega$[eckardt2015high]. In strongly correlated systems, such photon dressing significantly modifies the intermediate states involved in spin superexchange process [mentink2014ultrafast, mentink2015ultrafast]. As illustrated in Fig. 4a, an intermediate state with double occupancy transforms into photon-dressed states under periodic driving, effectively renormalizing the spin superexchange interactions. Consequently, the Floquet-engineered spin exchange interaction becomes [mentink2014ultrafast, mentink2015ultrafast]:
@@
where $\mathcal{J}_{m}(x)$ is the $m$th-order Bessel function of the first kind, describing contributions from processes involving $m$ photons with energy $\Omega$. Here, $E_{\rm int}$ represents the energy cost of creating a doubly occupied intermediate state relative to the singlet ground state. Crucially, the effective exchange interaction $J_{\alpha}^{\mathrm{eff}}$ depends on the projection $A_{0}\cdot\hat{\mathbf{e}}_{\alpha}$ of the pump field along the exchange pathway, allowing dynamic control over both the magnitude and the sign of the interaction by tuning the pump polarization, amplitude, and frequency [wang2021x].
@@
In contrast to the Hubbard model, where the intermediate-state energy is determined by $U$ alone [mentink2014ultrafast, mentink2015ultrafast], the EHM with on-site $U$ and nearest-neighbor $V$ results in an intermediate-state energy $E_{\mathrm{int}}\approx U-V$, setting the single-photon resonance at $\Omega\approx 9t_{h}$ in Eq. (4). Fig. 4b presents the simulated spin exchange for $A_{0}=0.4$ with $m$ truncated to $\pm 50$, where $J_{x}^{\rm eff}$ exhibits strong pump frequency dependence. Away from resonance, the Floquet-engineered effective $J_{x}^{\rm eff}$ recovers its equilibrium antiferromagnetic (AFM) exchange interaction $J\sim 4t_{h}^{2}/E_{\mathrm{int}}\sim 0.5t_{h}$. However, as $\Omega$ approaches resonance with $E_{\mathrm{int}}$, $J_{x}^{\rm eff}$ undergoes significant renormalization. Notably, when $\Omega$ slightly exceeds $E_{\rm int}$, the exchange interaction undergoes a sign reversal, transitioning from AFM to ferromagnetic (FM) (red-shaded region). This sign flip is crucial because AFM exchange favors spin-singlet pairing, while FM exchange supports spin-triplet pairing, thereby explaining the enhancement of $\Phi_{x}$ for $\Omega\gtrsim E_{\mathrm{int}}$. Moreover, the Floquet effect is polarization-selective: the exchange interaction along the $y$ direction ($J_{y}^{\rm eff}$) remains largely unaffected due to its orthogonality with the pump polarization, leading to the observed anisotropic pairing correlations.
@@
The resonance condition for Floquet-induced FM exchange is not confined to the narrow frequency window denoted by the red-shaded regime in Fig. 4b, but extends across a broader range of pump frequencies. This broadening arises from two primary effects. First, the above simplified Floquet analysis, which is based on atomic models, requires adjustments to incorporate band dispersion and quantum fluctuations in many-body systems. Second, the intermediate-state energy $E_{\mathrm{int}}$, which determines $J_{x}^{\rm eff}$ in Eq. (4), is not a fixed quantity in the doped EHM but varies across lattice sites in a given Fock state, depending on the local occupation configurations. As illustrated in the upper panels of Fig. 4c, configurations with isolated spin pairs or uniform neighbor densities yield $E_{\rm int}=U-V$, corresponding to a strong resonance near $\Omega\sim U-V=9t_{h}$ (solid lines in Fig. 4b). However, non-uniform occupations introduce a distribution of $E_{\rm int}$ values spanning from U4VU-4V to $U+2V$. At an average filling of $\langle n\rangle=0.5$, this variation extends the resonance window from $U+0.5V=7.5t_{h}$ to $U-2.5V=10.5t_{h}$, as indicated by translucence red and blue area in Fig. 4b. This asymmetric broadened resonance range around the center $\Omega=U-V$ aligns with the observed light-enhanced $p_{x}$ pairing instabilities in Fig. 3, reflecting the role of Floquet engineering in tuning pairing correlations.